A sailboat tracking against the wind moves as follows: 3.2km at 45°east of North, 4.5km at 50°west of North, 2.6km at 45°east of North. What is the net displacement of the entire motion?
1
Expert's answer
2020-08-27T10:25:27-0400

"a_x=3.2\cdot\sin45°-4.5\cdot\sin50°+2.6\cdot\sin45°=0.654(km)"


"a_y=3.2\cdot\cos45°+4.5\cdot\cos50°+2.6\cdot\cos45°=6.994(km)"


"a=\sqrt{0.654^2+6.994^2}=7.025(km)" Answer